The following data (and annotations) for March 2019 are for the work-in-process account of the first of Eyring Company’s four departments used in manufacturing its only product.

Work in Process-Department 1
Beginning balance (3,000 units, 40% complete)
Direct material $13,350
Conversion costs 5,600
Transferred to department 2: (23,000 units) (a)
Direct material (24,000 units) 96,000
Direct labor 77,300
Manufacturing overhead 34,700
Ending balance [ (b) units, 25% complete] (c)

Assuming that Erying uses the weighted average method and that materials are added at the beginning of the process and conversion costs are incurred evenly throughout, solve for the three missing numbers
Answer #1
Beginning work in process 3,000 completed & transferred 23,000
Started during the year 24,000 Ending work in process 4,000
units to be accounted 27,000 units accounted for 27,000
Equivalent units EUP EUP
units % DM DM % CC CC
completed & transferred 23,000 100% 23,000 100% 23,000
Ending work in process 4,000 100% 4,000 25% 1000
27,000 24,000
cost per unit
Materials CC total
Beginning work in process cost 13,350 5,600 18,950
cost added during the year 96,000 112000 208,000
total cost 109,350 117,600 226,950
Equivalent units 27,000 24,000
cost per unit 4.05 4.9 8.95
transferred to department 23,000 * 8.95 = 205850
Ending balance
materials 4,000 * 4.05 = 16200
Conversion 1000 * 4.9 = 4900 21100
total 226950
Answers)
A) 205850
B) 4,000 units
C) 21100